OpenStudy (anonymous):

Please help me solve this step-by-step. THANKS! 1.(D^2-16I)y=9.6(e^4x) + 30(e^x) 2.{2(D^2)-3D-2I}y=13-2(x^2) 3.(x^2)y''-(2x)y'+2y=(x^2)sinx

OpenStudy (anonymous):

1.\[(D^{2}-16I)y=9.6e ^{4x}+30e ^{x}\] 2.\[(2D ^{2}-3D-2I)y=13-2x ^{2}\] 3.\[x ^{2}y''-2xy'+2y=x ^{2}sinx\]